OpenStudy (ria23):

How do I convert y-3=-2(x+2)^2 to standard form?

OpenStudy (tkhunny):

Expand, expand, expand. (x+2)^2 = ???

OpenStudy (ria23):

... (x+2)(x+2)? Right?

OpenStudy (triciaal):

that's the start yes

OpenStudy (ria23):

So, y-3=-2(x+2)(x+2) Then I'm pretty sure yhu FOIL out the expansion next, right?

OpenStudy (triciaal):

then all that multiply by -2 next step add 3 to each side this gives y = standard form Ax + By = C so rearrange everything to get this form. whatever you do to one side of the equation also do to the other to keep the balance. When all is done then I will explain the shorter way to do the conversion. Just need to finish one approach at a time so you will not get confused.

OpenStudy (ria23):

y=-2x^2-8x-5?

OpenStudy (triciaal):

not quite -2* 4 =

OpenStudy (triciaal):

sorry you already added the 3

OpenStudy (ria23):

-8=2x^2-5

OpenStudy (triciaal):

sorrry standard ax^2 + bx = c

OpenStudy (triciaal):

sorry

OpenStudy (tkhunny):

No, don't "FOIL" anything. Just learn to multiply. (x+2)^2 = x^2 + 4x + 4

OpenStudy (triciaal):

standard for a line Ax + By = C standard for a quadratic ax^2 + bx + c = 0

OpenStudy (ria23):

But... @tkhunny ... That's what I got when... I FOILed it... O.o That's alright tricial c: So with it being in standard form, it would be -2x^2-8x-5=0... Right?

OpenStudy (ria23):

triciaal**

OpenStudy (triciaal):

yes

OpenStudy (tkhunny):

My point is, What are you going to do when you have to multiply these: (x^2 + x - 5)(2x^3 - 3x + 1) We'll need a new rule. Ff2O2fM2LIL2L - Perfect! Everyone memorize this! Seriously, forget "FOIL" and just learn to multiply.

OpenStudy (ria23):

O.O Woah... I'm only in pre-cal... I haven't even seen... Anything like that yet...

OpenStudy (tkhunny):

Again, that's my point. If you learn to multiply, and forget FOIL, that will not scare you when you see it. If all you know is FOIL, it will be a horrible stumbling block. Anyway, that's enough of that.

OpenStudy (ria23):

Alright. ^.^ Thanks for the heads up~

OpenStudy (triciaal):

@tkhunny what are you talking about? FOIL is just the code used to remember what to multiply

OpenStudy (tkhunny):

FOIL is a foolish code that has application ONLY for two binomials. It has no purpose on the Earth, unless you include confusing way, WAY more students than it helps.

OpenStudy (triciaal):

so stick to basics distribute multiply each term inside the parenthesis by what is outside. Personally never tell anyone to FOIL so I guess I agree

OpenStudy (ria23):

I grew up knowing that if the problem looks like that, then the easiest and fastest way to solve it would be to FOIL.

OpenStudy (ria23):

If there's more in the equation, then obviously yhu can't FOIL and yhu have to find a different method. But it's just what I've learned. It works, so I don't have any complaints.

OpenStudy (triciaal):

@Ria23 FOIL is just helping you to remember what to do. The operation like @tkhunny says is multiplication

OpenStudy (tkhunny):

Lot's of folks grew up thinking that. That's exactly why they NEVER learned how to multiply. I think it should be unceremoniously removed from all text books. Do you hear yourself talking? "have to find a different method" Yikes!!! No, you don't. If you learned to multiply, that method ALWAYS works. No need to start over when the silly method runs out of steam. FOIL will help you to remember how to multiply two binomials. It will not ever help you do anything else. Blech!! Throw the bum out!

OpenStudy (triciaal):

@Ria23 I think you also learn PEMDAS and don't know it is really a code to remember the order of operations.

OpenStudy (ria23):

Yea, I know PEMDAS. Is that what yhu guys are talking about when yhu say multiplying... Cuz I'm not really sure what yhu two mean. yhu say multiply and the first thing I think of is n x n= ?

OpenStudy (tkhunny):

And PEMDAS does many, many things and never runs out of applications. That is a mnemonic worth learning. Multiply these 25*43 = ??? How do you do that by hand? What numbers are multiplied by what numbers to produce the final result? Don't worry about the addition part. Just think about the manual multiplication process.

OpenStudy (triciaal):

I am saying that you know the codes used to remember the operation and don't know or understand the basics of the operation.

OpenStudy (ria23):

Yhu'd multiply 5 and 3, then 3 and 2... Then over the tens column.

OpenStudy (tkhunny):

Yup. Now, multiply (x+3)(y-2) EACTLY the same way. Match everything up just once. This is multiplication.

OpenStudy (ria23):

3 and -2, then 2 and x?

OpenStudy (ria23):

-2*

OpenStudy (triciaal):

25*43 = (20 + 5)(40 + 3)

OpenStudy (ria23):

O.o Oh god... That's.... Difficult to do just, off of the top of yhur head... 66 and 1700? O.O xD I feel like that's wrong....

OpenStudy (tkhunny):

@Ria23 See how odd that was? When you did the 2-digit numbers, you just naturally found all four pieces. When you did the binomials, you managed to find only two. I said EXACTLY the same way. Find all four combinations and you'll get it.

OpenStudy (ria23):

So... 20 and 40, and 5 and 3?

OpenStudy (tkhunny):

25*43 3*5 3*20 40*5 40*20 -- Four pieces. (x+3)(y-2) xy -3x 3y -6 -- Four pieces.

OpenStudy (ria23):

Wow that's... Incredibly confusing...

OpenStudy (tkhunny):

No, it's four pieces. You did it perfectly with two 2-digit numbers. You were not confused at all.

OpenStudy (tkhunny):

Anyway... y-3=-2(x+2)^2 y-3 = -2(x^2 + 4x + 4) y-3 = -2x^2 - 8x - 8 2x^2 + 8x + y - 3 + 8 = 0 2x^2 + 8x + y + 5 = 0

OpenStudy (triciaal):

(x+3)(y-2) xy -3x 3y -6 -- Four pieces. still four pieces minor error -3x should be -2x

OpenStudy (tkhunny):

Fair enough. It is better to get four CORRECT pieces. :-)